ELLOIT COMPANY Income Statement December 31, 2016 Cost of Goods ManufacturedThe following information is available for Elliot Company. January 1, 2016 $26,000 18,500 30,000 December 31, 2016 $30,000 22,200 21,000 2016 Raw materials inventory Work in process inventory Finished goods inventory Materials purchased Direct labor Manufacturing overhead Sales $170,000 230,000 180,000 800,000

Prepare an income statement through gross profit.
Answer #1

Direct material used = 26000+170000-30000 = 166000

Cost of goods manufactured = Beginning WIP+Direct material used+Direct labor+Manufacturing overhead-Ending WIP

= 18500+166000+230000+180000-22200

Cost of goods manufactured = 572300

Income statement

Sales 800000
Cost of goods sold
Beginning finished goods 30000
Add: Cost of goods manufactured 572300
Cost of goods available for sale 602300
Less: Ending finished goods -21000
Cost of goods sold 581300
Gross profit 218700